### v1

# LLM-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting

Abstract
Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios. Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios. Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting. However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series. To address this, we propose LLM-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment. Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts. Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information. Finally, the transformed time series from the LLMs are projected to obtain the forecasts. Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that LLM-Prompt is a powerful framework for time series forecasting.
1 Introduction
Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions [1]. Over the past five years, neural networks with diverse architectures—including linear models [2, 3], convolutional neural networks [4, 5], recurrent neural networks [6], and Transformers [7]—have been applied to time series forecasting. However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings [8, 9]. On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance [10, 11].
Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing. With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting [12, 13, 14]. However, due to the modality gap between LLMs and time series data, knowledge transfer is required. As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers [16]; b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering [9, 15]; c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e. a hybrid architecture of the above two strategies [17, 18].
Figure 1: (a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting. (b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts. (c) Multimodal-based LLMs: Use both text and time series as inputs but lack cross-modal alignment. (d) Our Method: Integrate multi-prompt information with time series through cross-modal alignment.
Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist. On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance [19, 20]. For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series [9]. Overall, a unified prompt paradigm remains absent. On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities [21, 22]. Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
To address the aforementioned technical challenges, this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed LLM-Prompt. The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture. Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data. Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation. Finally, the LLMs’ outputs are projected to obtain time series forecasting results. Experimental results demonstrate that LLM-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners. The main contributions of this study are summarized as follows:
  • We propose the LLM-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
  • We develop a synergistic mechanism integrating hard prompts and soft prompts, advancing the establishment of a unified prompt engineering paradigm for time series.
  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-fusion of temporal data and textual information, enhancing feature representation.
  • Extensive experiments on 7 benchmark datasets and 3 carbon emission datasets demonstrate that LLM-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
2 Related Work
Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures. Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures. iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms [23]. PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization [24]. Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series [25]. Other variants include Autoformer [26], which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer [27], which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting. DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency [2]. TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities [3]. CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years. SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions [28]. WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs [29]. RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently. Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting [6]. This method dynamically selects critical temporal features to capture long-term dependencies. However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios [32]. As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs. Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations. Prompt-based LLMs convert numerical time series into textual information. Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations. However, challenges remain in activating LLMs’ forecasting capabilities. Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts [17, 20]; LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data [8]; Time-LLM is limited to constructing hard prompts using only statistical features of temporal data [19]. Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap [30, 31].
3 Methodology
3.1 Framework Overview
As illustrated in Figure 2, LLM-Prompt comprises three key modules: dual-path prompting, cross-modal data fusion (including semantic space embedding and cross-modal alignment), and time series forecasting.
Figure 2: LLM-Prompt framework. First, hard prompts and soft prompts are constructed based on time series. Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment. Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
3.2 Dual-Path Prompting
Soft Prompts. Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection. The core idea is to dynamically select soft prompt vectors most relevant to the time series [17, 20]. First, initialize a prompt pool. Then, compute the similarity between the time series data and prompt keys in the pool, selecting the Top-k most relevant prompts. Notably, the prompt pool is adaptively updated during iterations.
Prompt Pool Construction. Given prompt keys $K\in{R^{P\times D}}$ and prompt pool $V\in{R^{P\times L\times D}}$, where P is the pool size, L is the soft prompt length, and D is the embedding dimension. The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs. Notably, the prompt pool is updated during network iterations. Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
Similarity Computation. Given the time series $X\in{R^{B\times N\times D}}$, where $B$ is the batch size, N is the variable dimension, and D is the embedding dimension. Compute the similarity matrix between X and K.
$S=Z{K^{T}}\in{R^{B\times N\times P}}$ (1)
Top-k Prompt Selection. Select the Top-k prompt key indices $I$ with the highest similarity. Retrieve the corresponding prompt vectors from the prompt pool using indices $I$ to construct soft prompts.
${\rm{softPromp}}{{\rm{t}}_{i}}=V[{I_{i}}]\in{R^{B\times N\times L\times D}}$ (2)
Figure 3: Hard prompt example
Hard Prompts. Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs. However, a well-considered unified paradigm remains absent. Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches [19, 21]. Experimental practice indicates this limits performance gains. To address this, we propose a unified hard prompt paradigm. As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features. Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities. Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
3.3 Cross-Modal Data Fusion
Semantic Space Embedding. This component primarily involves embedding time series data and hard prompts [21, 31, 22]. For soft prompts, they are generated directly in the embedding space and thus require no embedding.
Time Series. First, segment the time series into patches, then project them into the LLM’s embedding space to obtain $\hat{X}\in{R^{B\times N\times P\times D}}$, where P is the number of patches and $D$ is the embedding dimension. Given the pre-trained LLM’s vocabulary $E\in{R^{V\times D}}$, where $V$ is the vocabulary size, we realign time series data with the vocabulary through reprogramming, achieving textual representation of temporal data to enhance LLM comprehension. Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head. Then, multi-head self-attention produces the reprogramming temporal representations.
$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
Hard Prompt. Given the hard prompt $T$, map it to discrete tokens using the LLM’s pretrained tokenizer. Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer. During this process, both the tokenizer and LLM remain frozen with no parameter updates during training.
${\rm{hardPrompt}}={\rm{Embed}}({\rm{Tokenize}}(T))\in{R^{B\times N\times L\times D}}$ (4)
Cross-Modal Alignment. This module aims to achieve cross-modal alignment between soft-hard prompts and temporal data. Its core design resembles time series reprogramming in semantic space embedding. Specifically, cross-modal alignment is performed separately for soft/hard prompts and time series: (a) Dynamically align adaptive soft prompts with time series via a multi-head self-attention module, obtaining fused representation $Z_{1}$; (b) Process hard prompt text and time series using an identical architecture, obtaining fused representation $Z_{2}$. Finally, the joint representation $Z=[\tilde{X},Z_{1},Z_{2}]$ is formed through element-wise concatenation.
3.4 Time Series Forecasting
As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM. Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$. GPT-2 is used by default unless otherwise specified, and subsequent experiments also investigate the impact of different LLM architectures on forecasting performance [33].
4 Experiments
To validate the effectiveness of LLM-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting. Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs.
Baseline. We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM [19], TimeCMA [21]; (2) Transformer-based: Autoformer [26], Crossformer [35], iTransformer[23], PatchTST [36]; (3) CNN-based: TimeNet [5]; (4) RNN-based: Mamba[6]; (5) MLP-based: DLinear [2].
4.1 Long-term Forecasting
Setups. We conduct experiments on 9 datasets, including 6 classical benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, exchange-rate, weather) and 3 carbon emission datasets (munichCarbon, chinaCarbon, worldCarbon). The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others. Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
Results. The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that LLM-Prompt outperforms all baselines in most cases. LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations. Compared to other LLMs-based approaches, we observe that LLM-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA. This is attributed to our integration of both soft/hard prompts and cross-modal alignment in LLM-Prompt.
Table 1: Long-term forecasting results. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
4.2 Few-shot Forecasting
Setups. LLMs demonstrate remarkable few-shot learning capabilities in CV and NLP. We further investigate whether LLMs exhibit similar capabilities in forecasting tasks. Most settings follow the long-term forecasting configuration. This section evaluates model performance under scenarios with only 10% and 5% of training samples.
Results. The results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks. LLM-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
Table 2: Few-shot learning on 10% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.411 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
Table 3: Few-shot learning on 5% training data. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.491 0.469 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
ETTh2 0.388 0.408 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
ETTm1 0.418 0.411 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
ETTm2 0.289 0.333 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
Exchange 0.366 0.410 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
Weather 0.270 0.291 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
munichCarbon $0.419$ $0.472$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
chinaCarbon 0.720 0.610 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
worldCarbon 0.803 0.666 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
4.3 Zero-shot Forecasting
Table 4: Zero-shot learning on carbon emission datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE
$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
Setups. LLMs possess the potential for zero-shot time series forecasting. We further investigate the cross-domain zero-shot learning capabilities of LLMs. Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples. Most experimental settings follow the long-term forecasting configuration. We evaluate zero-shot learning performance on ETT-series and carbon emission datasets. Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
Results. Brief results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other deep learning approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting. In the $ETTh1\rightarrow ETTm1$ setting, LLM-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
Table 5: Zero-shot learning on ETT datasets. Red: the best, Blue: the second best.
Methods LLM-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$h1\rightarrow m1$ 96 0.723 0.553 $0.746$ 0.552 $0.772$ $0.565$ $0.766$ $0.573$ 0.713 $0.563$ $0.825$ $0.599$ $0.783$ $0.602$ $0.826$ $0.578$ $0.856$ $0.596$
192 $0.762$ 0.564 $0.832$ $0.587$ $0.807$ $0.581$ 0.749 0.568 0.737 $0.585$ $0.797$ $0.598$ $0.773$ $0.606$ $0.846$ $0.593$ $0.845$ $0.597$
336 0.755 0.557 $0.793$ 0.583 $0.807$ $0.586$ $0.843$ $0.589$ 0.757 $0.604$ $0.798$ $0.610$ $0.786$ $0.617$ $0.835$ $0.596$ $0.829$ $0.594$
720 0.765 0.584 $0.813$ 0.601 $0.819$ $0.601$ $0.865$ $0.605$ 0.800 $0.631$ $0.846$ $0.641$ $0.85$ $0.654$ $0.84$ $0.605$ $0.816$ $0.606$
Avg 0.751 0.564 $0.796$ 0.581 $0.801$ $0.583$ $0.806$ $0.584$ 0.752 $0.596$ $0.816$ $0.612$ $0.798$ $0.620$ $0.837$ $0.593$ $0.836$ $0.598$
$h2\rightarrow m2$ 96 0.213 0.301 0.221 $0.306$ $0.225$ $0.309$ $0.226$ $0.312$ $0.228$ $0.331$ $0.244$ $0.331$ $0.495$ $0.496$ $0.228$ $0.313$ 0.213 0.300
192 0.276 0.337 $0.279$ $0.337$ $0.286$ $0.341$ $0.287$ $0.343$ $0.300$ $0.379$ $0.298$ $0.358$ $0.638$ $0.562$ $0.287$ $0.344$ 0.273 0.334
336 0.330 0.365 0.335 0.370 $0.341$ $0.372$ $0.344$ $0.374$ $0.368$ $0.424$ $0.362$ $0.397$ $0.742$ $0.611$ $0.345$ $0.376$ $0.339$ $0.372$
720 0.432 0.420 $0.433$ 0.422 $0.438$ $0.426$ $0.443$ $0.431$ $0.514$ $0.507$ $0.470$ $0.456$ $1.017$ $0.741$ $0.442$ $0.427$ 0.429 0.420
Avg 0.313 0.356 $0.317$ $0.359$ $0.322$ $0.362$ $0.325$ $0.365$ $0.352$ $0.410$ $0.343$ $0.385$ $0.723$ $0.602$ $0.325$ $0.365$ 0.315 0.356
We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global). Results demonstrate that LLM-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
4.4 Ablation Study
We analyze the effectiveness of dual-path prompting and cross-modal alignment components. As shown in Table 7, ablations are conducted on multiple datasets (ETTm1, ETTh1, munichCarbon, chinaCarbon, worldCarbon). Removing either soft prompts, hard prompts, or cross-modal alignment degrades model performance. Prompt engineering is critical for large models to comprehend inputs: the absence of hard or soft prompts reduces performance by approximately 3.2%. The most severe impact occurs without cross-modal alignment, preventing the LLM from effectively understanding time series and causing a 3.7% performance drop.
Ablations on Hard Prompt. We further investigate how different types of information within hard prompts affect LLM-Prompt’s performance. As shown in Figure 3, we compare LLM-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration. As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃performance degradation.
Table 6: Ablations on different components. Red: the best.
Methods LLM-Prompt w/o Soft Prompt w/o Hard Prompt w/o Cross-Modal Alignment
Metric MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
Table 7: Ablations on different hard prompt configurations. Red: the best.
Methods LLM-Prompt w/o Historical Data w/o Statistical Data
Metric MSE MAE MSE MAE MSE MAE
ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
Ablations on Soft Prompt. Additionally, we investigate the impact of soft prompt pool size and soft prompt length on LLM-Prompt’s performance. As shown in Figure 5, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon. This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
Ablations on Parameter Sensitivity. We analyze the impact of intermediate projection dimension $d_{model}$ and multi-head attention head dimension $d_{ff}$ on model performance. As illustrated in Figure 5, varying parameter configurations have minimal impact on the ETTh1 dataset, possibly due to its sufficient sample size enabling effective model fitting. In contrast, for the chinaCarbon dataset, specific optimal configurations exist for both $d_{model}$ and $d_{ff}$, enabling optimal information flow within the network.
Figure 4: Ablations on different soft prompt parameters.
Figure 5: Ablations on different projection dimensions
5 Conclusion
This paper proposes LLM-Prompt to unlock the potential of LLMs in time series forecasting. The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks. To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features. Extensive experiments on multiple real-world datasets validate the effectiveness of LLM-Prompt in time series forecasting.
Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven. We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks. Ours future work will explore integrating Chain-of-Thought techniques to unlock LLMs’ predictive potential or developing foundation models tailored for forecasting tasks.

### v2

@@
-LLM-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting
+TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting
 Abstract
@@
 However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series.
-To address this, we propose LLM-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment.
+To address this, we propose TSF-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment.
 Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts.
@@
 Finally, the transformed time series from the LLMs are projected to obtain the forecasts.
-Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that LLM-Prompt is a powerful framework for time series forecasting.
-1 Introduction
-Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions [1].
-Over the past five years, neural networks with diverse architectures—including linear models [2, 3], convolutional neural networks [4, 5], recurrent neural networks [6], and Transformers [7]—have been applied to time series forecasting.
-However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings [8, 9].
-On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance [10, 11].
+Furthermore, we focus on applying time series forecasting to carbon emissions, aiming to provide a modest contribution to global carbon neutrality.
+Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that TSF-Prompt is a powerful framework for time series forecasting.
+Codehttps://github.com/SanMuGuo/Time-Prompt
+Introduction
+Time series forecasting, which utilizes past observations to predict future data, finds wide applications in domains such as electricity, weather, and carbon emissions (Wang et al.
+2024b).
+Over the past five years, neural networks with diverse architectures—including linear models (Zeng et al.
+2023; Wang et al.
+2024a), convolutional neural networks (Wang et al.
+2023; Wu et al.
+2023), recurrent neural networks (Gu and Dao 2023), and Transformers (Vaswani et al.
+2017)—have been applied to time series forecasting.
+However, existing models face two critical challenges: On one hand, they rely heavily on large-scale labeled samples for parameter optimization, often suffering from limited generalization capability in cross-domain scenarios, which results in performance barriers under few-shot or zero-shot forecasting settings (Gruver et al.
+2023; Liu et al.
+2024b).
+On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al.
+2024; Ni et al.
+2025).
 Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing.
-With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting [12, 13, 14].
+With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting (Shyalika et al.
+2024; Jin et al.
+2023b; Zhang et al.
+2024).
 However, due to the modality gap between LLMs and time series data, knowledge transfer is required.
-As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers [16]; b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering [9, 15]; c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e.
-a hybrid architecture of the above two strategies [17, 18].
+As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers (Zhou et al.
+2023); b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al.
+2024b; Xue and Salim 2023); c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e.
+a hybrid architecture of the above two strategies (Cao et al.
+2023; Jia et al.
+2024).
 Figure 1: (a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting.
@@
 Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist.
-On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance [19, 20].
-For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series [9].
+On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance (Jin et al.
+2023a; Pan et al.
+2024).
+For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series (Liu et al.
+2024b).
 Overall, a unified prompt paradigm remains absent.
-On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities [21, 22].
+On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities (Liu et al.
+2024a, 2025).
 Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
-To address the aforementioned technical challenges, this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed LLM-Prompt.
+To address the aforementioned technical challenges, inspired by recent research (Liu et al.
+2024a; Cao et al.
+2023; Jin et al.
+2023a), this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed TSF-Prompt.
 The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture.
@@
 Finally, the LLMs’ outputs are projected to obtain time series forecasting results.
-Experimental results demonstrate that LLM-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners.
+Experimental results demonstrate that TSF-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners.
+Furthermore, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals.
 The main contributions of this study are summarized as follows:
-  • We propose the LLM-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
+  • We propose the TSF-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
   • We develop a synergistic mechanism integrating hard prompts and soft prompts, advancing the establishment of a unified prompt engineering paradigm for time series.
-  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-fusion of temporal data and textual information, enhancing feature representation.
-  • Extensive experiments on 7 benchmark datasets and 3 carbon emission datasets demonstrate that LLM-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
-2 Related Work
+  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, enhancing feature representation.
+  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that TSF-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
+Related Work
 Deep learning has demonstrated superior performance in time series forecasting, primarily categorized into MLP-based, CNN-based, RNN-based, and Transformer-based architectures.
 Transformer-based models have emerged as the core research direction in recent years by capturing long-term dependencies and inter-variable correlations through self-attention mechanisms, achieving better predictive performance compared to other architectures.
-iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms [23].
-PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization [24].
-Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series [25].
-Other variants include Autoformer [26], which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer [27], which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
+iTransformer treats multivariate time series as independent variable sequences, capturing cross-variable dependencies via attention mechanisms (Liu et al.
+2023).
+PatchTST divides time series into fixed-length patches, enhancing local pattern recognition through patch-level tokenization (Nie et al.
+2023).
+Furthermore, addressing the challenge of non-stationary time series, Non-stationary Transformer proposes series stationarization and de-stationary attention to mitigate non-stationarity in raw time series (Liu et al.
+2022b).
+Other variants include Autoformer (Wu et al.
+2021), which integrates decomposition frameworks with autocorrelation mechanisms, and Fedformer (Zhou et al.
+2022), which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
+Figure 2: TSF-Prompt framework.
+First, hard prompts and soft prompts are constructed based on time series.
+Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment.
+Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
 Recent studies have challenged the effectiveness of Transformer-based models in time series forecasting.
-DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency [2].
-TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities [3].
+DLinear demonstrates that a simple linear regression in the original space achieves remarkable performance in both modeling and efficiency (Zeng et al.
+2023).
+TimeMixer introduces a multi-scale mixing mechanism, enhancing forecasting through feature fusion at different temporal granularities (Wang et al.
+2024a).
 CNN-based approaches, capturing local patterns via receptive fields, have also shown promising results in recent years.
-SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions [28].
-WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs [29].
+SCINet employs a hierarchical downsampling-convolution-interaction architecture to capture dynamic dependencies across temporal resolutions (Liu et al.
+2022a).
+WaveNet’s dilated causal convolutions significantly expand the receptive field, addressing long-term dependency issues in traditional CNNs (Van Den Oord et al.
+2016).
 RNN-based models, limited by vanishing gradients in long-term forecasting tasks, have seen fewer research efforts recently.
-Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting [6].
+Nevertheless, Mamba—based on state space models—has reestablished an effective paradigm for time series forecasting (Gu and Dao 2023).
 This method dynamically selects critical temporal features to capture long-term dependencies.
 However, current deep learning frameworks remain constrained by sample size, exhibiting suboptimal performance in few-shot or zero-shot scenarios.
-Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios [32].
+Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al.
+2024c).
 As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs.
@@
 However, challenges remain in activating LLMs’ forecasting capabilities.
-Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts [17, 20]; LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data [8]; Time-LLM is limited to constructing hard prompts using only statistical features of temporal data [19].
-Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap [30, 31].
-3 Methodology
-3.1 Framework Overview
-As illustrated in Figure 2, LLM-Prompt comprises three key modules: dual-path prompting, cross-modal data fusion (including semantic space embedding and cross-modal alignment), and time series forecasting.
-Figure 2: LLM-Prompt framework.
-First, hard prompts and soft prompts are constructed based on time series.
-Second, the time series and prompts are embedded into the LLM’s semantic space, followed by cross-modal alignment.
-Finally, the fused information is fed into the pretrained LLM and projected through output layers to generate forecasting results.
-3.2 Dual-Path Prompting
+Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts (Cao et al.
+2023; Pan et al.
+2024); LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data (Gruver et al.
+2023); Time-LLM is limited to constructing hard prompts using only statistical features of temporal data (Jin et al.
+2023a).
+Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap (Kowsher et al.
+2024; Hu et al.
+2025).
+Methodology
+Framework Overview
+As illustrated in Figure 2, TSF-Prompt comprises four components: dual-path prompting, semantic space embedding, cross-modal alignment, and time series forecasting.
+First, hard and soft prompts are constructed based on temporal data.
+Subsequently, semantic space embedding is performed through the pre-trained LLM and patch reprogramming module.
+Since soft prompts are inherently generated within this space, no additional embedding is required.
+Next, cross-modal alignment is executed.
+Finally, the fused information is fed into the pre-trained LLM and projected through output layers to generate forecasting results.
+Dual-Path Prompting
 Soft Prompts.
 Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection.
-The core idea is to dynamically select soft prompt vectors most relevant to the time series [17, 20].
+The core idea is to dynamically select soft prompt vectors most relevant to the time series (Cao et al.
+2023; Pan et al.
+2024).
 First, initialize a prompt pool.
@@
 The prompt pool can be initialized through multiple strategies: zero initialization, uniform distribution initialization, or initialized from the vocabulary of pretrained LLMs.
-Notably, the prompt pool is updated during network iterations.
+Notably, the prompt pool is updated during backpropagation.
 Subsequent experiments further investigate the impact of pool size and soft prompt length on model performance.
@@
 ${\rm{softPromp}}{{\rm{t}}_{i}}=V[{I_{i}}]\in{R^{B\times N\times L\times D}}$ (2)
-Figure 3: Hard prompt example
+Figure 3: Hard prompt example on chinaCarbon dataset.
 Hard Prompts.
@@
 However, a well-considered unified paradigm remains absent.
-Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches [19, 21].
+Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches (Jin et al.
+2023a; Liu et al.
+2024a).
 Experimental practice indicates this limits performance gains.
@@
 Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
-3.3 Cross-Modal Data Fusion
+Table 1: Long-term forecasting results. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
+ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
+ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
+ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
+Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
+Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
+munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
+chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
+worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
+Table 2: Few-shot learning on 10% training data. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.461 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
+ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
+ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
+ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
+Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
+Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
+munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
+chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
+worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
+Cross-Modal Data Fusion
 Semantic Space Embedding.
-This component primarily involves embedding time series data and hard prompts [21, 31, 22].
+This component primarily involves embedding time series data and hard prompts (Liu et al.
+2024a; Hu et al.
+2025; Liu et al.
+2025).
 For soft prompts, they are generated directly in the embedding space and thus require no embedding.
@@
 Finally, the joint representation $Z=[\tilde{X},Z_{1},Z_{2}]$ is formed through element-wise concatenation.
-3.4 Time Series Forecasting
+Time Series Forecasting
 As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM.
 Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$.
-GPT-2 is used by default unless otherwise specified, and subsequent experiments also investigate the impact of different LLM architectures on forecasting performance [33].
-4 Experiments
-To validate the effectiveness of LLM-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting.
+GPT-2 is used by default unless otherwise specified (Radford et al.
+2019).
+Experiments
+To validate the effectiveness of TSF-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting.
 Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs.
+Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
+All experiments are conducted on V100 GPUs.
+Additional details (e.g., datasets, SOTA methods, full results) are provided in the appendix.
 Baseline.
-We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM [19], TimeCMA [21]; (2) Transformer-based: Autoformer [26], Crossformer [35], iTransformer[23], PatchTST [36]; (3) CNN-based: TimeNet [5]; (4) RNN-based: Mamba[6]; (5) MLP-based: DLinear [2].
-4.1 Long-term Forecasting
+We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al.
+2023a), TimeCMA (Liu et al.
+2024a); (2) Transformer-based: Autoformer (Wu et al.
+2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al.
+2023), PatchTST (Nie et al.
+2022); (3) CNN-based: TimeNet (Wu et al.
+2023); (4) RNN-based: Mamba(Gu and Dao 2023); (5) MLP-based: DLinear (Zeng et al.
+2023).
+Long-term Forecasting
 Setups.
@@
 The input sequence length $T$ is fixed at 96, and we adopt four different forecasting horizons: $H\in\{60,90,120,150\}$ for chinaCarbon and worldCarbon, and $H\in\{96,192,336,720\}$ for the others.
-Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
 Results.
-The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that LLM-Prompt outperforms all baselines in most cases.
+The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that TSF-Prompt outperforms all baselines in most cases.
 LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations.
-Compared to other LLMs-based approaches, we observe that LLM-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA.
-This is attributed to our integration of both soft/hard prompts and cross-modal alignment in LLM-Prompt.
-Table 1: Long-term forecasting results. Red: the best, Blue: the second best.
-Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
-ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
-ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
-ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
-Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
-Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
-munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
-chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
-worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
-4.2 Few-shot Forecasting
-Setups.
-LLMs demonstrate remarkable few-shot learning capabilities in CV and NLP.
-We further investigate whether LLMs exhibit similar capabilities in forecasting tasks.
-Most settings follow the long-term forecasting configuration.
-This section evaluates model performance under scenarios with only 10% and 5% of training samples.
-Results.
-The results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks.
-LLM-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
-Table 2: Few-shot learning on 10% training data. Red: the best, Blue: the second best.
-Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.411 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
-ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
-ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
-ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
-Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
-Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
-munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
-chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
-worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
-Table 3: Few-shot learning on 5% training data. Red: the best, Blue: the second best.
-Methods LLM-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
+Compared to other LLMs-based approaches, we observe that TSF-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA.
+This is attributed to our integration of both soft/hard prompts and cross-modal alignment in TSF-Prompt.
+Table 3: Few-shot learning on 5% training data. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 worldCarbon 0.803 0.666 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
-4.3 Zero-shot Forecasting
-Table 4: Zero-shot learning on carbon emission datasets. Red: the best, Blue: the second best.
-Methods LLM-Prompt Time-LLM iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE
-$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
-$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
-$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
-$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
-$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
-$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
-Setups.
-LLMs possess the potential for zero-shot time series forecasting.
-We further investigate the cross-domain zero-shot learning capabilities of LLMs.
-Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples.
-Most experimental settings follow the long-term forecasting configuration.
-We evaluate zero-shot learning performance on ETT-series and carbon emission datasets.
-Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
-Results.
-Brief results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other deep learning approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting.
-In the $ETTh1\rightarrow ETTm1$ setting, LLM-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
-Table 5: Zero-shot learning on ETT datasets. Red: the best, Blue: the second best.
-Methods LLM-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
+Table 4: Zero-shot learning on ETT datasets. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 Avg 0.313 0.356 $0.317$ $0.359$ $0.322$ $0.362$ $0.325$ $0.365$ $0.352$ $0.410$ $0.343$ $0.385$ $0.723$ $0.602$ $0.325$ $0.365$ 0.315 0.356
+Few-shot Forecasting
+Setups.
+LLMs demonstrate remarkable few-shot learning capabilities in CV and NLP.
+We further investigate whether LLMs exhibit similar capabilities in forecasting tasks.
+Most settings follow the long-term forecasting configuration.
+This section evaluates model performance under scenarios with only 10% and 5% of training samples.
+Results.
+The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks.
+TSF-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
+Zero-shot Forecasting
+Setups.
+LLMs possess the potential for zero-shot time series forecasting.
+We further investigate the cross-domain zero-shot learning capabilities of LLMs.
+Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples.
+Most experimental settings follow the long-term forecasting configuration.
+We evaluate zero-shot learning performance on ETT-series and carbon emission datasets.
+Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
+Results.
+Results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting.
+In the $ETTh1\rightarrow ETTm1$ setting, TSF-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
 We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global).
-Results demonstrate that LLM-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
-4.4 Ablation Study
+Results demonstrate that TSF-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
+Table 5: Zero-shot learning on carbon emission datasets. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM iTransformer PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE
+$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
+$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
+$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
+$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
+$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
+$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
+Table 6: Ablations on different components. Bold: the best.
+Methods TSF-Prompt w/o Soft Prompt w/o Hard Prompt w/o Alignment
+Metric MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
+ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
+munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
+chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
+worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
+Table 7: Ablations on different hard prompt configurations.
+Methods TSF-Prompt w/o Historical Data w/o Statistical Data
+Metric MSE MAE MSE MAE MSE MAE
+ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
+chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
+Figure 4: Ablations on different soft prompt parameters.
+Figure 5: Ablations on different projection dimensions
+Ablation Study
 We analyze the effectiveness of dual-path prompting and cross-modal alignment components.
@@
 Ablations on Hard Prompt.
-We further investigate how different types of information within hard prompts affect LLM-Prompt’s performance.
-As shown in Figure 3, we compare LLM-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration.
-As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃performance degradation.
-Table 6: Ablations on different components. Red: the best.
-Methods LLM-Prompt w/o Soft Prompt w/o Hard Prompt w/o Cross-Modal Alignment
-Metric MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
-ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
-munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
-chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
-worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
-Table 7: Ablations on different hard prompt configurations. Red: the best.
-Methods LLM-Prompt w/o Historical Data w/o Statistical Data
-Metric MSE MAE MSE MAE MSE MAE
-ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
-chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
+We further investigate how different types of information within hard prompts affect TSF-Prompt’s performance.
+As shown in Figure 3, we compare TSF-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration.
+As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃% performance degradation.
 Ablations on Soft Prompt.
-Additionally, we investigate the impact of soft prompt pool size and soft prompt length on LLM-Prompt’s performance.
+Additionally, we investigate the impact of soft prompt pool size and soft prompt length on TSF-Prompt’s performance.
 As shown in Figure 4, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon.
@@
 In contrast, for the chinaCarbon dataset, specific optimal configurations exist for both $d_{model}$ and $d_{ff}$, enabling optimal information flow within the network.
-Figure 4: Ablations on different soft prompt parameters.
-Figure 5: Ablations on different projection dimensions
-5 Conclusion
-This paper proposes LLM-Prompt to unlock the potential of LLMs in time series forecasting.
+Conclusion
+This paper proposes TSF-Prompt to unlock the potential of LLMs in time series forecasting.
 The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks.
 To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features.
-Extensive experiments on multiple real-world datasets validate the effectiveness of LLM-Prompt in time series forecasting.
+Extensive experiments on multiple real-world datasets validate the effectiveness of TSF-Prompt in time series forecasting.
 Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven.
 We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks.
-Ours future work will explore integrating Chain-of-Thought techniques to unlock LLMs’ predictive potential or developing foundation models tailored for forecasting tasks.
+Our future work will explore developing foundation models specifically tailored for forecasting tasks, or further demonstrate the predictive capabilities of LLMs through techniques such as chain-of-thought reasoning, fine-tuning, and reinforcement learning.

### v3

@@
-TSF-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting
+Time-Prompt: Integrated Heterogeneous Prompts for Unlocking LLMs in Time Series Forecasting
 Abstract
 Time series forecasting aims to model temporal dependencies among variables for future state inference, holding significant importance and widespread applications in real-world scenarios.
-Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting and data-scarce scenarios.
-Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting.
-However, we find existing LLM-based methods still have shortcomings: (1) the absence of a unified paradigm for textual prompt formulation and (2) the neglect of modality discrepancies between textual prompts and time series.
-To address this, we propose TSF-Prompt, an LLM-based time series forecasting framework integrating multi-prompt information and cross-modal semantic alignment.
-Specifically, we first construct a unified textual prompt paradigm containing learnable soft prompts and textualized hard prompts.
-Second, to enhance LLMs’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve cross-modal fusion of temporal and textual information.
-Finally, the transformed time series from the LLMs are projected to obtain the forecasts.
-Furthermore, we focus on applying time series forecasting to carbon emissions, aiming to provide a modest contribution to global carbon neutrality.
-Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that TSF-Prompt is a powerful framework for time series forecasting.
-Codehttps://github.com/SanMuGuo/Time-Prompt
+Although deep learning-based methods have achieved remarkable progress, they still exhibit suboptimal performance in long-term forecasting.
+Recent research demonstrates that large language models (LLMs) achieve promising performance in time series forecasting, but this progress is still met with skepticism about whether LLMs are truly useful for this task.
+To address this, we propose Time-Prompt, a framework for activating LLMs for time series forecasting.
+Specifically, we first construct a unified prompt paradigm with learnable soft prompts to guide the LLM’s behavior and textualized hard prompts to enhance the time series representations.
+Second, to enhance LLM’ comprehensive understanding of the forecasting task, we design a semantic space embedding and cross-modal alignment module to achieve fusion of temporal and textual data.
+Finally, we efficiently fine-tune the LLM’s parameters using time series data.
+Furthermore, we focus on carbon emissions, aiming to provide a modest contribution to global carbon neutrality.
+Comprehensive evaluations on 6 public datasets and 3 carbon emission datasets demonstrate that Time-Prompt is a powerful framework for time series forecasting.
 Introduction
@@
 2023; Liu et al.
-2024b).
+2024a).
 On the other hand, existing methods struggle to capture long-term temporal dependencies, resulting in suboptimal long-term forecasting performance (Chen et al.
@@
 Pretrained foundation models, such as Large Language Models (LLMs), have demonstrated strong few-shot and zero-shot learning capabilities in computer vision and natural language processing.
-With the technological advancements driven by LLMs, increasing research efforts focus on leveraging their potential for time series forecasting (Shyalika et al.
+As pre-trained models advance, research is increasingly split between harnessing LLMs for time series forecasting (Shyalika et al.
 2024; Jin et al.
-2023b; Zhang et al.
+2023; Zhang et al.
+2024) and training time series foundation models (Goswami et al.
+2024; Ansari et al.
+2024; Wiliński et al.
 2024).
-However, due to the modality gap between LLMs and time series data, knowledge transfer is required.
-As shown in Figure 1, the most common current approaches are: a) Time series-based LLMs: achieving modality alignment from time series to the semantic space of LLMs via learnable embedding layers (Zhou et al.
-2023); b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al.
-2024b; Xue and Salim 2023); c) Multimodal-based LLMs: using both time series and textual information as multimodal inputs, i.e.
-a hybrid architecture of the above two strategies (Cao et al.
-2023; Jia et al.
+Considering the high computational cost of foundation models, our work focuses on the former.
+As shown in Figure 1, the most common LLMs-based approaches are: (a) Time series-based LLMs: Projecting the time series into the LLM’s semantic space via an embedding layer (Zhou et al.
+2023); (b) Prompt-based LLMs: converting temporal data into textual modality through prompt engineering (Liu et al.
+2024a); (c) Multimodal-based LLMs: Processing of time series and text separately (Zhou et al.
+2025); (d) Multimodal-based LLMs: Feeding unified multimodal information into the LLMs (Zhao, Chen, and Sun 2025).
+Figure 1: a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting.
+(b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts.
+(c) Multimodal-based LLMs: Time series are processed via an encoder, and text via LLMs, separately.
+(d) Multimodal-based LLMs: Time series and text information are processed by LLMs in a unified manner.
+Currently, research primarily focuses on two main types (c) and (d).
+For example, TimeCMA first encodes the multimodal data, where the time series is processed by an encoder and the text by a frozen LLM.
+Subsequently, cross-modal alignment is performed, and finally, the prediction result is obtained through a decoder (Liu et al.
+2025a).
+In contrast, Time-LLM processes both time series and text uniformly through a frozen LLM, which then projects the target value (Jin et al.
 2024).
-Figure 1: (a) Time series-based LLMs: Lack textual information, failing to leverage LLMs’ potential for time series forecasting.
-(b) Prompt-based LLMs: Struggle to capture fine-grained temporal patterns from textual prompts.
-(c) Multimodal-based LLMs: Use both text and time series as inputs but lack cross-modal alignment.
-(d) Our Method: Integrate multi-prompt information with time series through cross-modal alignment.
-Current state-of-the-art approaches predominantly employ multimodal-based LLMs, yet notable limitations persist.
-On one hand, existing prompt engineering strategies utilize either hard prompts or soft prompts exclusively, failing to explore their synergy for further enhancing LLM performance (Jin et al.
-2023a; Pan et al.
-2024).
-For hard prompts, they either solely utilize statistical information of time series or directly textualize raw time series (Liu et al.
-2024b).
-Overall, a unified prompt paradigm remains absent.
-On the other hand, most works employ simple concatenation for multimodal data fusion, inadequately addressing the semantic gap and feature redundancy across different modalities (Liu et al.
-2024a, 2025).
-Notably, these technical limitations directly lead to significant shortcomings in long-term dependency modeling and noise robustness.
-To address the aforementioned technical challenges, inspired by recent research (Liu et al.
-2024a; Cao et al.
-2023; Jin et al.
-2023a), this paper proposes an LLM-empowered time series forecasting framework based on a collaborative mechanism of hard and soft prompts, termed TSF-Prompt.
-The core innovations of this framework lie in two aspects: a) constructing a unified paradigm for soft-hard prompt synergy and b) establishing a cross-modal information fusion network architecture.
-Specifically, we first design a dual-path prompt template integrating hard prompts (including temporal statistical features and raw numerical values) and adaptive soft prompts (optimized via gradient backpropagation) based on existing time series data.
-Then, we build a cross-modal data fusion module, including a semantic space embedding module and a cross-modal alignment module, to embed raw time series data and dual-path prompt texts into the LLMs’ dimension followed by modality adaptation.
-Finally, the LLMs’ outputs are projected to obtain time series forecasting results.
-Experimental results demonstrate that TSF-Prompt is effective for time series forecasting and verifies LLMs as capable few-shot and zero-shot time series learners.
-Furthermore, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals.
+Although current methods have achieved state-of-the-art performance, there remains some skepticism—do LLM truly benefit time series forecasting?
+We believe that with appropriate guidance, the time series forecasting potential of LLMs can be activated.
+This is because, for LLMs, textual reasoning is fundamentally a form of discrete-value reasoning.
+Viewed from this perspective, it is quite similar to the task of time series forecasting, with the main distinction being that time series consists of continuous values.
+Furthermore, LLMs have exhibited certain mathematical reasoning abilities, which also indicates a capacity for handling continuous values.
+To unlock the potential of LLMs for time series forecasting, inspired by recent research (Liu et al.
+2025a; Cao et al.
+2024; Pan et al.
+2024; Jin et al.
+2024), we propose Time-Prompt, an LLM-empowered framework that activates the potential of LLMs for time series forecasting through time series reprogramming, dual-path prompting, cross-modal alignment, and LoRA.
+We design a dual-path prompting framework that activates the LLM’s specific patterns with soft prompts and enhances its time series representations with hard prompts.
+We also reprogram the time series to endow it with a text-like representation.
+Furthermore, we construct a cross-modal alignment module to eliminate the modality gap between text and time, and we efficiently fine-tune the LLM with LoRA to enhance its ability to process continuous values.
+In addition, given that carbon emissions have become a globally critical issue, we conduct a cross-scale study on carbon emission forecasting—spanning global, national, and urban levels—to provide technical support for the strategic requirements of global carbon neutrality goals.
 The main contributions of this study are summarized as follows:
-  • We propose the TSF-Prompt framework, incorporating a soft-hard prompt synergy mechanism and cross-modal data fusion modules to address limitations in long-term dependency modeling and cross-domain generalization.
-  • We develop a synergistic mechanism integrating hard prompts and soft prompts, advancing the establishment of a unified prompt engineering paradigm for time series.
-  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, enhancing feature representation.
-  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that TSF-Prompt achieves superior performance, particularly in few-shot and zero-shot scenarios.
+  • We propose Time-Prompt, a framework that activates the potential of LLMs in time series forecasting through strategies such as dual-path prompting, cross-modal alignment, and efficient fine-tuning.
+  • We design a dual-path prompting strategy, with learnable soft prompts guiding the LLM’s behavior and hard prompts extracting enhanced time series representations.
+  • We construct a semantic space embedding module and a cross-modal alignment module to enable cross-modal fusion of temporal data and textual information, eliminating the modality gap.
+  • Extensive experiments on 6 benchmark datasets and 3 carbon emission datasets demonstrate that Time-Prompt achieves superior performance.
 Related Work
@@
 2022), which employs frequency-domain enhanced decomposition strategies, both significantly improving long-term forecasting accuracy.
-Figure 2: TSF-Prompt framework.
+Figure 2: Time-Prompt framework.
 First, hard prompts and soft prompts are constructed based on time series.
@@
 Recently, leveraging the cross-domain transfer capabilities of LLMs, they have demonstrated exceptional performance in time series forecasting tasks, significantly outperforming traditional methods, especially in few-shot scenarios (Liu et al.
-2024c).
-As shown in Figure 1, the most prevalent current approaches include: a) Time series-based LLMs; b) Prompt-based LLMs; c) Multimodal-based LLMs.
-Time series-based LLMs achieve cross-modal alignment from time series to the LLM semantic space via learnable embedding layers, encoding temporal data into LLM-readable representations.
-Prompt-based LLMs convert numerical time series into textual information.
-Current research focuses on Multimodal-based LLMs, which perform cross-modal semantic alignment between temporal and textual data to overcome inherent limitations of single-modality representations.
-However, challenges remain in activating LLMs’ forecasting capabilities.
-Specifically, prompt engineering lacks a unified paradigm: TEMPO and S2IP-LLM solely employ learnable soft prompts, neglecting the importance of hard prompts (Cao et al.
-2023; Pan et al.
-2024); LLM-Time directly textualizes temporal data as hard prompts for LLMs, resulting in feature redundancy due to high overlap with raw temporal data (Gruver et al.
-2023); Time-LLM is limited to constructing hard prompts using only statistical features of temporal data (Jin et al.
-2023a).
-Furthermore, existing multimodal alignment methods rely on feature concatenation, failing to effectively bridge the modality gap (Kowsher et al.
-2024; Hu et al.
+2024b).
+As shown in Figure 1, the most prevalent current approaches include: (a) Time series-based LLMs; (b) Prompt-based LLMs; (c) Multimodal-based LLMs; (d) Multimodal-based LLMs.
+Time series–based LLMs project time series into the semantic space of LLMs via an embedding layer, and subsequently map the embedding vectors to obtain prediction values (Zhou et al.
+2023).
+In contrast, Prompt-based LLMs construct the time series as a text modality (Xue and Salim 2023).
+Lstprompt is a typical Prompt-based method that proposes a prompt construction technique, guiding the LLM to make predictions via a Chain-of-Thought approach (Liu et al.
+2024a).
+However, single-modality time series forecasting often suffers from limited representational capacity, leading to suboptimal prediction performance.
+Therefore, many studies attempt to integrate Type (a) and Type (b) approaches, leveraging multimodal information to overcome the inherent limitations of single-modality methods.
+Early efforts predominantly fall into Type (d), which uniformly feeds time series and textual inputs into LLMs to obtain improved feature representations.
+GPT4MTS, a foundational work in this category, embeds time series and simple text prompts into the semantic space, from which the large model derives a joint representation (Jia et al.
+2024).
+Many subsequent works build upon this framework.
+Time-LLM enhances model comprehension of time series by reprogramming the input time series (Jin et al.
+2024).
+Meanwhile, S²IP-LLM and TEMPO employ learnable soft prompts to mark specific temporal patterns, thereby more effectively activating the large model (Cao et al.
+2024; Pan et al.
+2024).
+Subsequently, with growing skepticism about whether LLMs are truly useful for time series (Tan et al.
+2024), some studies shift away from feeding time series into LLMs for representation and instead adopt dedicated encoders—this approach corresponds to Type (c).
+TimeCMA encodes the time series with an encoder and the text with an LLM, followed by cross-modal alignment, and finally generates the prediction via a unified decoder (Liu et al.
+2025a).
+In a related approach, another study further enhances the textual information using reinforcement learning (Su et al.
 2025).
+The problem with Type (c) is that the LLM only plays an auxiliary role, contributing little to time series forecasting.
+The issue with Type (d), however, is the LLM’s inability to handle continuous values like those in time series.
+Nevertheless, recent research continues to emerge for both Type (c) (Zhou et al.
+2025; Su et al.
+2025) and Type (d) (Chen et al.
+2025; Zhang et al.
+2025), in an effort to address the widespread skepticism.
+Our method belongs to the Type (d).
+LLMs possess powerful text-based reasoning capabilities.
+Fundamentally, their primary distinction from time series lies in the nature of data: discrete versus continuous values.
+Therefore, time series forecasting and textual reasoning exhibit significant similarity.
+Moreover, LLMs have also demonstrated strong reasoning abilities in mathematics.
+So we believe that with appropriate guidance, LLMs can enhance the performance of time series forecasting.
+Our method mainly contributes in this area.
+On the LLM side, we guide the LLM’s latent patterns with soft prompts and achieve efficient fine-tuning through LoRA.
+For the time series aspect, we reprogram the time series data and incorporate cross-modal alignment.
 Methodology
 Framework Overview
-As illustrated in Figure 2, TSF-Prompt comprises four components: dual-path prompting, semantic space embedding, cross-modal alignment, and time series forecasting.
-First, hard and soft prompts are constructed based on temporal data.
-Subsequently, semantic space embedding is performed through the pre-trained LLM and patch reprogramming module.
-Since soft prompts are inherently generated within this space, no additional embedding is required.
-Next, cross-modal alignment is executed.
-Finally, the fused information is fed into the pre-trained LLM and projected through output layers to generate forecasting results.
+The overall framework of Time-Prompt is illustrated in Figure 2.
+First, we construct a dual-path prompting framework for the time series forecasting task.
+This framework incorporates both hard prompts (including temporal statistical features and historical data) and learnable soft prompts.
+Our goal is to guide the LLM’s internal behavioral patterns via soft prompts and enhance time series feature representation through hard prompts.
+Second, we embed both the time series and text into the LLM’s semantic space via semantic space embedding module to facilitate better comprehension by the LLM.
+Subsequently, to address the modality gap between text and time series, we employ cross-modal alignment.
+Finally, the global and local features of the time series are extracted and projected to obtain the predicted values.
+Here, the LLM is fine-tuned using LoRA to adapt it from discrete-value reasoning to continuous-value prediction.
 Dual-Path Prompting
@@
 Soft prompts include prompt pool construction, similarity computation, and Top-k prompt selection.
-The core idea is to dynamically select soft prompt vectors most relevant to the time series (Cao et al.
-2023; Pan et al.
-2024).
+The core idea is to dynamically select soft prompt vectors most relevant to the time series.
 First, initialize a prompt pool.
@@
 Hard Prompts.
-Hard prompts provide a direct yet effective approach to activate forecasting capabilities in LLMs.
-However, a well-considered unified paradigm remains absent.
-Based on our observation, existing works either solely textualize temporal data or construct text using only statistical features of time series, failing to integrate both approaches (Jin et al.
-2023a; Liu et al.
-2024a).
-Experimental practice indicates this limits performance gains.
-To address this, we propose a unified hard prompt paradigm.
-As illustrated in Figure 3, our paradigm comprises four key components: (1) Dataset Description, (2) Task Description, (3) Historical Data, and (4) Statistical Features.
-Dataset description provides background context for LLMs, while task description, historical data, and statistical features collectively enrich temporal information to enhance the LLMs’ forecasting capabilities.
-Subsequent experiments demonstrate that both auxiliary information [(1), (2), (3)] and temporal data [(4)] are all effective.
+Hard prompts provide a direct and effective way to activate the LLM’s time series forecasting capability, as no modality gap exists between the LLM and the prompts.
+Hard prompts can also be viewed as enhanced representations of the time series.
+As illustrated in Figure 3, our paradigm comprises three key components: (1) Instruction, (2) Statistical Characteristics, and (3) Historical Data.
+The Instruction provides the LLM with the task context, the Statistical characteristics correctly guide the LLM’s patterns, and the Historical Data is the direct textualization of the time series data.
 Table 1: Long-term forecasting results. Bold: the best, Underline: the second best.
@@
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.433 0.438 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
-ETTh2 0.369 0.398 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
-ETTm1 0.381 0.399 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
-ETTm2 0.283 0.329 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
-Exchange 0.355 0.400 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
-Weather 0.254 0.279 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
-munichCarbon $0.412$ $0.465$ $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ 0.415 0.469 $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
-chinaCarbon 0.584 0.540 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
-worldCarbon $0.646$ $0.563$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
+ETTh1 0.430 0.439 $0.449$ 0.445 $0.449$ $0.447$ $0.520$ $0.492$ $0.460$ $0.455$ $0.460$ $0.457$ $0.477$ $0.475$ $0.549$ $0.530$ $0.464$ $0.455$ 0.446 $0.447$
+ETTh2 0.368 0.397 0.380 0.403 $0.399$ $0.416$ $0.435$ $0.447$ $0.408$ $0.421$ $0.564$ $0.519$ $0.446$ $0.457$ $1.920$ $1.073$ $0.383$ $0.406$ $0.382$ $0.411$
+ETTm1 0.380 0.398 $0.421$ $0.413$ $0.400$ $0.410$ $0.466$ $0.450$ $0.408$ $0.416$ $0.404$ $0.408$ $0.53$ $0.496$ $0.697$ $0.605$ $0.407$ $0.412$ 0.390 0.405
+ETTm2 0.281 0.326 $0.291$ $0.335$ $0.294$ $0.336$ $0.331$ $0.365$ $0.299$ 0.333 $0.354$ $0.401$ $0.326$ $0.364$ $1.725$ $0.813$ $0.291$ $0.334$ 0.293 $0.334$
+Exchange 0.331 0.390 $0.361$ 0.403 $0.448$ $0.451$ $0.761$ $0.584$ $0.425$ $0.447$ 0.340 $0.414$ $0.843$ $0.596$ $0.757$ $0.645$ $0.364$ $0.405$ $0.391$ $0.417$
+Weather 0.252 0.277 $0.266$ $0.285$ $0.257$ $0.284$ $0.288$ $0.314$ $0.258$ $0.285$ $0.265$ $0.317$ $0.337$ $0.379$ $0.259$ $0.320$ $0.261$ $0.281$ 0.255 0.278
+munichCarbon 0.411 0.464 $0.411$ $0.465$ $0.433$ $0.486$ $0.485$ $0.518$ $0.415$ $0.469$ $0.394$ $0.456$ $0.611$ $0.599$ 0.358 0.437 $0.412$ $0.465$ $0.430$ $0.477$
+chinaCarbon 0.571 0.532 $0.627$ $0.557$ $0.659$ $0.586$ $0.660$ $0.592$ $0.665$ $0.588$ $0.689$ $0.642$ $0.642$ $0.608$ $1.177$ $0.834$ $0.667$ $0.579$ 0.604 0.546
+worldCarbon $0.617$ $0.544$ $0.649$ $0.565$ 0.546 0.518 $0.687$ $0.614$ $0.751$ $0.639$ $0.754$ $0.698$ $0.653$ $0.603$ $0.619$ 0.539 $0.735$ $0.616$ 0.615 $0.547$
 Table 2: Few-shot learning on 10% training data. Bold: the best, Underline: the second best.
@@
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.461 0.452 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
-ETTh2 0.383 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
-ETTm1 0.411 0.408 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
-ETTm2 0.288 0.332 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
-Exchange 0.339 0.397 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
-Weather $0.270$ $0.289$ 0.269 0.288 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
-munichCarbon 0.412 0.465 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
-chinaCarbon 0.659 0.577 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
-worldCarbon $0.732$ $0.622$ $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ 0.725 0.609
+ETTh1 0.452 0.445 0.477 $0.460$ $0.549$ $0.504$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ $0.510$ $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.477$ 0.455
+ETTh2 0.381 0.404 $0.396$ $0.414$ $0.416$ $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ 0.391 0.412
+ETTm1 0.394 0.402 $0.424$ $0.411$ $0.425$ $0.427$ $0.511$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ 0.414 0.406
+ETTm2 0.284 0.327 $0.297$ $0.339$ $0.302$ $0.342$ $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ 0.290 0.333
+Exchange 0.337 0.395 $0.368$ 0.407 $0.423$ $0.448$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.136$ $0.411$ $0.442$ $0.373$ $0.411$
+Weather 0.265 0.286 $0.269$ $0.288$ 0.255 0.282 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.339$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.273$ $0.289$
+munichCarbon 0.406 0.459 $0.426$ $0.477$ $0.434$ $0.486$ $0.513$ $0.538$ $0.445$ $0.491$ $0.413$ $0.475$ $0.605$ $0.609$ 0.402 0.474 $0.443$ $0.491$ $0.423$ $0.474$
+chinaCarbon 0.653 0.575 $0.714$ $0.606$ $0.690$ $0.596$ $0.785$ $0.651$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.043$ $0.833$ $0.792$ $0.645$ 0.665 0.580
+worldCarbon 0.705 0.601 $0.802$ $0.669$ 0.720 0.607 $0.950$ $0.733$ $0.959$ $0.740$ $0.845$ $0.699$ $0.781$ $0.659$ $0.966$ $0.782$ $0.937$ $0.727$ $0.725$ $0.609$
 Cross-Modal Data Fusion
 Semantic Space Embedding.
-This component primarily involves embedding time series data and hard prompts (Liu et al.
-2024a; Hu et al.
-2025; Liu et al.
-2025).
+This component primarily involves embedding the time series and prompts into the LLM’s semantic space to enhance the LLM’s understanding of the variables.
 For soft prompts, they are generated directly in the embedding space and thus require no embedding.
@@
 Specifically, learnable projection matrices generate query matrices $Q=\hat{X}{W^{Q}}$, key matrices $K=E{W^{K}}$, and value matrices $V=E{W^{V}}$ for each attention head.
-Then, multi-head self-attention produces the reprogramming temporal representations.
-$\tilde{X}={\rm{MHSA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
+Then, multi-head cross-attention produces the reprogramming temporal representations.
+$\tilde{X}={\rm{MHCA}}(Q,K,V)\in{R^{B\times N\times P\times D}}$ (3)
 Hard Prompt.
@@
 Then, convert these discrete tokens into continuous vector representations via the LLM’s input embedding layer.
-During this process, both the tokenizer and LLM remain frozen with no parameter updates during training.
+During this process, the LLM’s tokenizer remain frozen with no parameter updates during training.
 ${\rm{hardPrompt}}={\rm{Embed}}({\rm{Tokenize}}(T))\in{R^{B\times N\times L\times D}}$ (4)
 Cross-Modal Alignment.
-This module aims to achieve cross-modal alignment between soft-hard prompts and temporal data.
-Its core design resembles time series reprogramming in semantic space embedding.
-Specifically, cross-modal alignment is performed separately for soft/hard prompts and time series: (a) Dynamically align adaptive soft prompts with time series via a multi-head self-attention module, obtaining fused representation $Z_{1}$; (b) Process hard prompt text and time series using an identical architecture, obtaining fused representation $Z_{2}$.
-Finally, the joint representation $Z=[\tilde{X},Z_{1},Z_{2}]$ is formed through element-wise concatenation.
+This module is designed to achieve cross-modal fusion among the time series data, soft prompts, and hard prompts to eliminate the modality gap.
+Many existing works simply concatenate these variables, which can lead to severe feature redundancy when the number of prompt tokens is large (Jin et al.
+2024).
+Our specific approach is as follows: (a) We use a multi-head cross-attention mechanism to align the soft prompt with the time series, obtaining a fused representation $Z_{1}$.
+Similarly, the hard prompt is aligned with the time series to obtain $Z_{2}$.
+(b) We then fuse the original time series and the fused representations $Z_{1}$ and $Z_{2}$ via a gating mechanism, yielding the final representation $Z$.
+$Z=g_{0}\odot\tilde{X}+g_{1}\odot Z_{1}+g_{2}\odot Z_{2}$ (5)
+Table 3: Few-shot learning on 5% training data. Bold: the best, Underline: the second best.
+Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.464 0.454 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
+ETTh2 0.387 0.406 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
+ETTm1 0.400 0.404 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
+ETTm2 0.285 0.329 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
+Exchange 0.358 0.403 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
+Weather 0.268 0.288 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
+munichCarbon $0.416$ $0.469$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
+chinaCarbon 0.654 0.577 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
+worldCarbon 0.738 0.619 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
+Table 4: Ablation Experiment
+Methods Time-Prompt w/o SP w/o HP w/o CMA w/o LoRA Qwen 3 0.6B
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.430 0.439 $0.434$ $0.441$ $0.437$ $0.442$ $0.434$ $0.442$ $0.435$ $0.440$ $0.717$ $0.579$
+ETTm1 0.380 0.398 $0.383$ $0.399$ $0.384$ $0.400$ $0.386$ $0.402$ $0.386$ $0.401$ $0.719$ $0.564$
+munichCarbon 0.411 0.532 $0.413$ $0.466$ $0.414$ $0.467$ $0.414$ $0.467$ $0.416$ $0.468$ $0.652$ $0.588$
+chinaCarbon 0.571 0.532 $0.581$ $0.540$ $0.584$ $0.540$ $0.587$ $0.544$ $0.583$ $0.538$ $0.802$ $0.652$
+worldCarbon 0.617 0.544 $0.636$ $0.558$ $0.633$ $0.556$ $0.633$ $0.556$ $0.630$ $0.552$ $0.869$ $0.713$
 Time Series Forecasting
-As shown in Figure 2, the joint representation $Z$ is fed into the frozen LLM.
-Subsequently, its output is passed to a projection layer to obtain the final prediction $Y$.
-GPT-2 is used by default unless otherwise specified (Radford et al.
-2019).
+Many existing works rely solely on frozen LLMs.
+Given the high dimensionality of the LLM’s semantic space, they often resort to direct truncation to prevent the curse of dimensionality.
+This approach, however, is detrimental to activating the LLM’s potential for time series forecasting.
+In this module, we first employ LoRA to efficiently fine-tune the LLM, enhancing its ability to comprehend continuous values.
+Second, to address the excessively high semantic dimensionality, we project its representations into a lower-dimensional space, extract global features, and then jointly project these global features together with local features to produce the final prediction $Y$.
+$F_{\text{local}}$ $=\text{Linear}(Z),$ (6)
+$F_{\text{global}}$ $=\text{Mean}(F_{\text{local}}),$ (7)
+$F_{\text{fused}}$ $=\text{Linear}([F_{\text{local}};F_{\text{global}}]),$ (8)
+$Y$ $=\text{Head}(\text{Mean}(F_{\text{fused}})).$ (9)
 Experiments
-To validate the effectiveness of TSF-Prompt, we conduct extensive experiments across various time series forecasting tasks, including long-term forecasting, few-shot forecasting, and zero-shot forecasting.
-Additionally, we investigate the positive effects of soft prompts and hard prompts in enhancing the forecasting capabilities of LLMs.
-Evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE).
-All experiments are conducted on V100 GPUs.
-Additional details (e.g., datasets, SOTA methods, full results) are provided in the appendix.
+Dataset.
+We utilize six widely-used public datasets covering multiple domains, including electricity, finance, and weather, as well as three carbon emission datasets with different spatial scales.
+Among them, the Munich carbon emission dataset was collected by one of our co-authors.
+We conduct extensive experiments across these nine datasets to demonstrate the superiority of Time-Prompt, including tasks such as long-term forecasting and few-shot forecasting.
+Zero-shot forecasting is not considered in this work, as we think LLM-based frameworks currently lack strong generalization capability in such settings.
 Baseline.
 We select 9 representative SOTAs from 5 categories: (1) LLMs-based: Time-LLM (Jin et al.
-2023a), TimeCMA (Liu et al.
-2024a); (2) Transformer-based: Autoformer (Wu et al.
+2024), TimeCMA (Liu et al.
+2025a); (2) Transformer-based: Autoformer (Wu et al.
 2021), Crossformer (Zhang and Yan 2023), iTransformer(Liu et al.
@@
 2023).
+Implementation.
+Since the vast majority of LLM-based time series forecasting methods adopt GPT-2, we also use GPT-2 by default in this work for fair comparison (Radford et al.
+2019).
+The evaluation metrics are Mean Squared Error (MSE) and Mean Absolute Error (MAE).
+Additional details (e.g., datasets, baseline, full results) are provided in the appendix.
 Long-term Forecasting
@@
 Results.
-The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that TSF-Prompt outperforms all baselines in most cases.
-LLM-based methods surpass deep learning and linear models in long-term forecasting, aligning with one of this work’s motivations.
-Compared to other LLMs-based approaches, we observe that TSF-Prompt achieves a reduction of 3.5%/1.53% in average MSE/MAE over Time-LLM and 4.25%/3.07% over Time-CMA.
-This is attributed to our integration of both soft/hard prompts and cross-modal alignment in TSF-Prompt.
-Table 3: Few-shot learning on 5% training data. Bold: the best, Underline: the second best.
-Methods TSF-Prompt Time-LLM TimeCMA Mamba TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.491 0.469 $0.527$ $0.491$ $0.583$ $0.521$ $0.716$ $0.572$ $0.847$ $0.632$ $0.632$ $0.548$ 0.510 $0.499$ $0.583$ $0.531$ $0.701$ $0.570$ $0.518$ 0.479
-ETTh2 0.388 0.408 $0.409$ $0.424$ 0.417 $0.429$ $0.427$ $0.448$ $0.465$ $0.459$ $0.435$ $0.458$ $0.427$ $0.446$ $0.800$ $0.642$ $0.445$ $0.447$ $0.418$ 0.424
-ETTm1 0.418 0.411 $0.425$ $0.413$ 0.435 0.431 $0.510$ $0.467$ $0.474$ $0.454$ $0.496$ $0.467$ $0.623$ $0.532$ $0.427$ $0.439$ $0.491$ $0.456$ $0.433$ $0.419$
-ETTm2 0.289 0.333 $0.294$ $0.337$ 0.304 0.343 $0.318$ $0.359$ $0.311$ $0.352$ $0.382$ $0.431$ $0.311$ $0.355$ $0.592$ $0.527$ $0.304$ $0.345$ $0.303$ $0.348$
-Exchange 0.366 0.410 $0.389$ $0.417$ $0.407$ $0.436$ $0.677$ $0.554$ $0.439$ $0.459$ 0.368 $0.431$ $0.471$ $0.483$ $2.043$ $1.135$ $0.411$ $0.442$ $0.370$ 0.411
-Weather 0.270 0.291 $0.273$ $0.293$ 0.264 0.290 $0.293$ $0.322$ $0.276$ $0.301$ $0.275$ $0.338$ $0.348$ $0.380$ $0.343$ $0.397$ $0.281$ $0.295$ $0.280$ $0.296$
-munichCarbon $0.419$ $0.472$ $0.437$ $0.487$ $0.435$ $0.486$ $0.513$ $0.537$ $0.445$ $0.491$ 0.402 0.469 $0.604$ $0.609$ 0.341 0.435 $0.443$ $0.491$ $0.426$ $0.476$
-chinaCarbon 0.720 0.610 $0.743$ $0.623$ 0.721 0.612 $0.785$ $0.652$ $0.821$ $0.655$ $0.742$ $0.632$ $0.766$ $0.652$ $1.042$ $0.832$ $0.792$ $0.645$ $0.732$ $0.613$
-worldCarbon 0.803 0.666 $0.859$ $0.703$ 0.745 0.631 $0.950$ $0.733$ $0.959$ $0.740$ $0.846$ $0.699$ $0.821$ $0.682$ $0.965$ $0.782$ $0.937$ $0.727$ $0.824$ $0.671$
-Table 4: Zero-shot learning on ETT datasets. Bold: the best, Underline: the second best.
-Methods TSF-Prompt Time-LLM TimeCMA TimesNet DLinear Autoformer Crossformer iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-$h1\rightarrow m1$ 96 0.723 0.553 $0.746$ 0.552 $0.772$ $0.565$ $0.766$ $0.573$ 0.713 $0.563$ $0.825$ $0.599$ $0.783$ $0.602$ $0.826$ $0.578$ $0.856$ $0.596$
-192 $0.762$ 0.564 $0.832$ $0.587$ $0.807$ $0.581$ 0.749 0.568 0.737 $0.585$ $0.797$ $0.598$ $0.773$ $0.606$ $0.846$ $0.593$ $0.845$ $0.597$
-336 0.755 0.557 $0.793$ 0.583 $0.807$ $0.586$ $0.843$ $0.589$ 0.757 $0.604$ $0.798$ $0.610$ $0.786$ $0.617$ $0.835$ $0.596$ $0.829$ $0.594$
-720 0.765 0.584 $0.813$ 0.601 $0.819$ $0.601$ $0.865$ $0.605$ 0.800 $0.631$ $0.846$ $0.641$ $0.85$ $0.654$ $0.84$ $0.605$ $0.816$ $0.606$
-Avg 0.751 0.564 $0.796$ 0.581 $0.801$ $0.583$ $0.806$ $0.584$ 0.752 $0.596$ $0.816$ $0.612$ $0.798$ $0.620$ $0.837$ $0.593$ $0.836$ $0.598$
-$h2\rightarrow m2$ 96 0.213 0.301 0.221 $0.306$ $0.225$ $0.309$ $0.226$ $0.312$ $0.228$ $0.331$ $0.244$ $0.331$ $0.495$ $0.496$ $0.228$ $0.313$ 0.213 0.300
-192 0.276 0.337 $0.279$ $0.337$ $0.286$ $0.341$ $0.287$ $0.343$ $0.300$ $0.379$ $0.298$ $0.358$ $0.638$ $0.562$ $0.287$ $0.344$ 0.273 0.334
-336 0.330 0.365 0.335 0.370 $0.341$ $0.372$ $0.344$ $0.374$ $0.368$ $0.424$ $0.362$ $0.397$ $0.742$ $0.611$ $0.345$ $0.376$ $0.339$ $0.372$
-720 0.432 0.420 $0.433$ 0.422 $0.438$ $0.426$ $0.443$ $0.431$ $0.514$ $0.507$ $0.470$ $0.456$ $1.017$ $0.741$ $0.442$ $0.427$ 0.429 0.420
-Avg 0.313 0.356 $0.317$ $0.359$ $0.322$ $0.362$ $0.325$ $0.365$ $0.352$ $0.410$ $0.343$ $0.385$ $0.723$ $0.602$ $0.325$ $0.365$ 0.315 0.356
+The results summarized in Table 1, averaged over different forecasting horizons, demonstrate that Time-Prompt outperforms all baselines in most cases.
+Time-Prompt achieves the best performance in 14 out of 18 benchmark settings.
+Overall, LLM-based methods demonstrate superior performance compared to conventional deep learning approaches.
+Specifically, Time-Prompt reduces MSE and MAE by 5.5% and 2.7% over Time-LLM, by 6.3% and 4.2% over TimeCMA, and by 4.3% and 2.4% compared to PatchTST—the strongest deep learning baseline.
 Few-shot Forecasting
@@
 The brief results summarized in Tables 2 and 3, averaged over different forecasting horizons, demonstrate that LLMs-based methods significantly outperform deep learning models, demonstrating that LLMs retain few-shot learning capabilities in forecasting tasks.
-TSF-Prompt outperforms other LLMs-based approaches, which we attribute to the synergistic integration of soft/hard prompts better activating LLMs’ few-shot learning capabilities.
-Zero-shot Forecasting
-Setups.
-LLMs possess the potential for zero-shot time series forecasting.
-We further investigate the cross-domain zero-shot learning capabilities of LLMs.
-Specifically, models are trained on one dataset and tested on another dataset with entirely unseen samples.
-Most experimental settings follow the long-term forecasting configuration.
-We evaluate zero-shot learning performance on ETT-series and carbon emission datasets.
-Due to differing variable dimensions in carbon emission datasets, only channel-independent models are evaluated.
-Results.
-Results in Tables 4 and 5 show that LLMs-based methods and PatchTST outperform other approaches in cross-domain scenarios, demonstrating LLMs’ established zero-shot reasoning capability for time series forecasting.
-In the $ETTh1\rightarrow ETTm1$ setting, TSF-Prompt achieves reductions of 5.64% and 2.8% in MSE/MAE over Time-LLM and 10.1% and 5.64% over PatchTST.
-We further comprehensively investigate LLMs’ zero-shot learning capabilities for carbon emission forecasting through cross-domain experiments across three spatial scales (city, country, and global).
-Results demonstrate that TSF-Prompt frameworks serve as effective zero-shot predictors for carbon emissions, providing a technical foundation for achieving global carbon neutrality.
-Table 5: Zero-shot learning on carbon emission datasets. Bold: the best, Underline: the second best.
-Methods TSF-Prompt Time-LLM iTransformer PatchTST
-Metric MSE MAE MSE MAE MSE MAE MSE MAE
-$munich\rightarrow china$ 0.687 0.591 0.697 0.596 $0.731$ $0.607$ $0.797$ $0.633$
-$munich\rightarrow world$ 0.897 0.705 0.934 0.719 $1.036$ $0.753$ $0.987$ $0.737$
-$china\rightarrow munich$ 0.432 0.482 $0.449$ $0.495$ 0.446 $0.492$ $0.447$ 0.490
-$china\rightarrow world$ 0.791 0.656 $0.818$ $0.664$ $0.876$ $0.694$ 0.740 0.613
-$world\rightarrow munich$ 0.445 0.487 0.445 $0.492$ $0.448$ 0.491 $0.475$ $0.502$
-$world\rightarrow china$ 0.631 $0.562$ $0.644$ 0.561 $0.689$ $0.587$ 0.616 0.545
-Table 6: Ablations on different components. Bold: the best.
-Methods TSF-Prompt w/o Soft Prompt w/o Hard Prompt w/o Alignment
-Metric MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1 0.433 0.438 $0.444$ $0.442$ $0.443$ $0.441$ $0.443$ $0.441$
-ETTm1 0.381 0.399 $0.402$ $0.408$ $0.403$ $0.409$ $0.410$ $0.409$
-munichCarbon 0.412 0.465 $0.422$ $0.476$ $0.423$ $0.476$ $0.422$ $0.475$
-chinaCarbon 0.584 0.540 $0.608$ $0.549$ $0.609$ $0.551$ $0.608$ $0.550$
-worldCarbon 0.646 0.563 $0.660$ $0.572$ $0.658$ $0.575$ $0.665$ $0.577$
-Table 7: Ablations on different hard prompt configurations.
-Methods TSF-Prompt w/o Historical Data w/o Statistical Data
-Metric MSE MAE MSE MAE MSE MAE
-ETTh1 - 96 0.391 0.406 $0.389$ $0.406$ $0.389$ $0.406$
-chinaCarbon - 96 0.472 0.477 $0.495$ $0.490$ $0.499$ $0.491$
-Figure 4: Ablations on different soft prompt parameters.
-Figure 5: Ablations on different projection dimensions
-Ablation Study
-We analyze the effectiveness of dual-path prompting and cross-modal alignment components.
-As shown in Table 6, ablations are conducted on multiple datasets (ETTm1, ETTh1, munichCarbon, chinaCarbon, worldCarbon).
-Removing either soft prompts, hard prompts, or cross-modal alignment degrades model performance.
-Prompt engineering is critical for large models to comprehend inputs: the absence of hard or soft prompts reduces performance by approximately 3.2%.
-The most severe impact occurs without cross-modal alignment, preventing the LLM from effectively understanding time series and causing a 3.7% performance drop.
-Ablations on Hard Prompt.
-We further investigate how different types of information within hard prompts affect TSF-Prompt’s performance.
-As shown in Figure 3, we compare TSF-Prompt’s performance under three scenarios: (1) only auxiliary information, (2) only temporal data, and (3) full integration.
-As presented in Table 7, varying hard prompt types have minimal impact on the ETTh1 dataset but significantly affect the chinaCarbon dataset, where missing information causes 5̃% performance degradation.
-Ablations on Soft Prompt.
-Additionally, we investigate the impact of soft prompt pool size and soft prompt length on TSF-Prompt’s performance.
-As shown in Figure 4, varying parameters have minimal effects on the ETTh1 dataset, while extreme parameter values (too large or too small) cause performance degradation on chinaCarbon.
-This may be attributed to information redundancy from oversized parameters or insufficient soft prompt information from undersized parameters.
-Ablations on Parameter Sensitivity.
-We analyze the impact of intermediate projection dimension $d_{model}$ and multi-head attention head dimension $d_{ff}$ on model performance.
-As illustrated in Figure 4, varying parameter configurations have minimal impact on the ETTh1 dataset, possibly due to its sufficient sample size enabling effective model fitting.
-In contrast, for the chinaCarbon dataset, specific optimal configurations exist for both $d_{model}$ and $d_{ff}$, enabling optimal information flow within the network.
+The superiority of Time-Prompt is even more pronounced in few-shot scenarios compared to long-term forecasting.
+Time-Prompt achieves the best performance in 30 out of 36 benchmark settings Under the 10% training data setting, Time-Prompt reduces MSE and MAE by 7.1% and 4.3% over Time-LLM, and by 8.0% and 5.5% over TimeCMA.
+Under the more challenging 5% training data setting, it further achieves 8.8% and 5.7% MSE/MAE reductions compared to Time-LLM, and 7.9% and 5.5% reductions compared to TimeCMA.
+Table 5: Modal Analysis.
+soft prompt 3 5 10 15
+MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.376 0.402 0.376 0.402 0.374 0.401 0.377 0.402
+munichCarbon 0.238 0.352 0.239 0.352 0.237 0.351 0.236 0.351
+chinaCarbon 0.473 0.472 0.462 0.469 0.458 0.469 0.467 0.473
+hard prompt 3 5 10 15
+MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.378 0.405 0.376 0.402 0.376 0.402 0.375 0.400
+munichCarbon 0.236 0.350 0.239 0.352 0.242 0.354 0.236 0.350
+chinaCarbon 0.464 0.474 0.462 0.469 0.466 0.469 0.460 0.467
+d_model 32 64 128 256
+MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.376 0.402 0.378 0.402 0.378 0.402 0.381 0.403
+munichCarbon 0.239 0.352 0.236 0.349 0.241 0.352 0.238 0.350
+chinaCarbon 0.462 0.469 0.467 0.473 0.470 0.477 0.455 0.468
+d_ff 32 64 128 256
+MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.381 0.405 0.380 0.406 0.376 0.402 0.380 0.400
+munichCarbon 0.239 0.352 0.243 0.354 0.239 0.352 0.239 0.351
+chinaCarbon 0.482 0.483 0.460 0.469 0.462 0.469 0.464 0.471
+Ablation Experiment
+As shown in Table 4, we conduct an ablation study on multiple datasets (ETTh1, ETTm1, munichCarbon, chinaCarbon, worldCarbon) to analyze the effects of soft prompts (SP), hard prompts (HP), cross-modal alignment (CMA), and LoRA.
+All experiments follow the long-term forecasting setting, using four different prediction horizons, and the reported results are averaged across these settings.
+The results indicate that all these modules activate the LLM’s time series forecasting capabilities to varying degrees.
+Compared to the prompting engineering, LoRA and cross-modal alignment have a more significant impact on the LLM’s forecasting ability.
+Additionally, we replace GPT-2 with the Qwen-3 0.6B model, which lead to a significant performance drop.
+We hypothesize that this is because larger-parameter models struggle to effectively handle continuous-valued time-series data.
+This observation aligns with findings in the time-series foundation model literature: most published time-series foundation models are relatively small, typically in the few-hundred-million parameter range or even smaller (Team 2025).
+This also indicates that in future research, we should focus more on the design of model architecture and the tuning of model parameters, rather than solely on prompt engineering.
+Furthermore, greater attention should be paid to foundation models and small-scale LLMs (Liu et al.
+2025b; Shi et al.
+2025).
+Modal Analysis
+We experiment with the settings for soft prompt length, hard prompt length, $d_{model}$, and $d_{ff}$ on multiple datasets (ETTh1, munichCarbon, chinaCarbon).
+The hard prompt length refers to the number of tokens retained after the hard prompt is encoded by the LLM.
+Due to the relatively large number of tokens, we truncate it to mitigate the curse of dimensionality.
+This approach is supported by existing literature, which indicates that the key information is predominantly located in the latter part (Liu et al.
+2025a).
+Here, $d_{model}$ represents the intermediate dimension of the patch reprogram module, while $d_{ff}$ is the intermediate dimension of the projection module.
+As shown in the Table 5, the soft prompt length should neither be too short nor too long.
+Since soft prompts are designed to guide the LLM’s behavior, an excessively long sequence of tokens is unnecessary.
+In contrast, hard prompts are constructed by textualizing time-series data, so retaining more tokens is generally better, as truncation inevitably leads to information loss.
+However, considering the trade-off between efficiency and performance, we opt to keep only the most informative segment.
+For $d_{ff}$ and $d_{model}$, we employ a grid search to select the optimal parameter combination.
+Ultimately, for all datasets, we default to $d_{ff}=128$ and $d_{model}=32$.
 Conclusion
-This paper proposes TSF-Prompt to unlock the potential of LLMs in time series forecasting.
-The method incorporates a dual-path prompting mechanism, where the synergistic integration of dynamic selection of learnable soft prompts and semantic mapping of textualized hard prompts enhances LLMs’ comprehension of forecasting tasks.
-To further improve performance, a cross-modal data fusion architecture is designed through multi-head cross-attention modules, achieving alignment of textual and temporal features.
-Extensive experiments on multiple real-world datasets validate the effectiveness of TSF-Prompt in time series forecasting.
-Despite initial progress, we observe that sole reliance on prompt engineering exhibits significant limitations in effectively activating LLMs’ forecasting capabilities—or, more critically, the intrinsic forecasting capability of LLMs remains unproven.
-We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to numerical tasks.
-Our future work will explore developing foundation models specifically tailored for forecasting tasks, or further demonstrate the predictive capabilities of LLMs through techniques such as chain-of-thought reasoning, fine-tuning, and reinforcement learning.
+This paper proposes Time-Prompt to unlock the potential of LLMs in time series forecasting.
+This method employs a dual-path prompting mechanism.
+It utilizes soft prompts to guide the reasoning patterns of the LLM and enhances representations with hard prompts and time series reprogramming.
+Subsequently, it bridges the modality gap between time series and text through semantic space embedding and cross-modal alignment.
+Finally, the LLM’s capability for processing continuous values is enhanced through LoRA fine-tuning.
+Extensive experiments on multiple real-world datasets validate the effectiveness of Time-Prompt in time series forecasting.
+Despite preliminary advancements, we find that the time series forecasting capabilities of LLMs remain inadequate.
+Whether the discrete reasoning abilities of language models can be effective for time series forecasting is still an open question.
+We argue that this fundamental limitation stems from the text-centric training paradigm of LLMs, which inherently struggles to adapt to forecasting tasks.
+Moreover, according to the experiments in this paper, we find that small-scale LLMs are more effective than LLMs, and fine-tuning is more effective than prompt engineering in activating predictive capabilities—especially in few-shot scenarios.
+So our future work will be dedicated to developing time series foundation models or exploring the time-series forecasting capabilities of small-scale LLMs.